Evaluate the ∫∫ z dS where S is the upper half of a sphere of radius 4.
1
Expert's answer
2020-04-21T16:30:59-0400

We use the spherical coordinates

"x=r \sin \varphi \cos \theta,\ \ \ y=r \sin \varphi \sin \theta ,\ \ \ z= \cos \varphi"

Since "r=4" , then the parameterization for this sphere

"\vec{r}(\theta, \varphi)=4 \sin \varphi \cos \theta \vec{\imath}+4 \sin \varphi \sin \theta \vec{\jmath}+4 \cos \varphi \vec{k}"

Since we are working on the upper half of the sphere here are the limits on the parameters.

"0 \leq \theta \leq 2 \pi \quad 0 \leq \varphi \leq \frac{\pi}{2}"

Since

"\begin{array}{l}
\vec{r}_{\theta}(\theta, \varphi)=-2 \sin \varphi \sin \theta \vec{i}+2 \sin \varphi \cos \theta \vec{j} \\
\vec{r}_{\varphi}(\theta, \varphi)=2 \cos \varphi \cos \theta \vec{i}+2 \cos \varphi \sin \theta \vec{j}-2 \sin \varphi \vec{k}
\end{array}"

Then

"\begin{aligned}
\vec{r}_{\theta} \times \vec{r}_{\varphi} &=\left|\begin{array}{cc}
\vec{i} & \vec{j} & \vec{k} \\
-4\sin \varphi \sin \theta & 4 \sin \varphi \cos \theta & 0 \\
4 \cos \varphi \cos \theta & 4 \cos \varphi \sin \theta & -4\sin \varphi
\end{array}\right| \\
&=-16 \sin ^{2} \varphi \cos \theta \vec{i}-16 \sin \varphi \cos \varphi \sin ^{2} \theta \vec{k}-16 \sin ^{2} \varphi \sin \theta \vec{j}\\
&\ \ \ \ \ \ -16 \sin \varphi \cos \varphi \cos ^{2} \theta \vec{k} \\
&=-16 \sin ^{2} \varphi \cos \theta \vec{i}-16 \sin ^{2} \varphi \sin \theta \vec{j}\\
&\ \ \ \ \ -16 \sin \varphi \cos \varphi\left(\sin ^{2} \theta+\cos ^{2} \theta\right) \vec{k} \\
&=-16 \sin ^{2} \varphi \cos \theta \vec{i}-16 \sin ^{2} \varphi \sin \theta \vec{j} -16 \sin \varphi \cos \varphi \vec{k}
\end{aligned}"

Hence the magnitude

"\begin{aligned}
\left\|\vec{r}_{\theta} \times \vec{r}_{\varphi}\right\| &=\sqrt{16^2 \sin ^{4} \varphi \cos ^{2} \theta+16^2 \sin ^{4} \varphi \sin ^{2} \theta+16 ^2\sin ^{2} \varphi \cos ^{2} \varphi} \\
&=\sqrt{16^2 \sin ^{4} \varphi\left(\cos ^{2} \theta+\sin ^{2} \theta\right)+16^2 \sin ^{2} \varphi \cos ^{2} \varphi} \\
&=\sqrt{16^2 \sin ^{2} \varphi\left(\sin ^{2} \varphi+\cos ^{2} \varphi\right)}\\
&=16\sin \varphi 
\end{aligned}"

The surface integral is 

"\begin{aligned}
\iint\limits_{S}{{z\,dS}} &= \iint\limits_{D}{{4\cos \varphi \left( {16\sin \varphi } \right)\,dA}}\\
 & = \int_{{\,0}}^{{\,2\pi }}{{\int_{{\,0}}^{{\,\frac{\pi }{2}}}{{32\sin \left( {2\varphi } \right)\,d\varphi }}\,d\theta }}\\ & = \int_{{\,0}}^{{\,2\pi }}{{ {\left( { - 16\cos \left( {2\varphi } \right)} \right)} \bigg|_0^{\frac{\pi }{2}}\,d\theta }}\\ & = \int_{{\,0}}^{{\,2\pi }}{{16\,d\theta }}\\ & = 32\pi 
\end{aligned}"